OpenStudy (anonymous):

Dan confirms the authenticity of an amethyst stone by a few tests. During one of the tests, he finds the angle of the refracted ray of light coming out of the amethyst into the air as 28°. If the refractive index of amethyst is 1.55, what is the incident angle? a- 152 degree b-26 degree c-47 degree d-71 degree

OpenStudy (aaronq):

do you know what equation to use?

OpenStudy (anonymous):

no:/

OpenStudy (anonymous):

do you?

OpenStudy (aaronq):

yep, so use: \[\eta _{1}\sin _{1} \Theta=\eta _{2}\sin \theta _{2}\]

OpenStudy (aaronq):

n is the refractive index

OpenStudy (aaronq):

n for air is 1

OpenStudy (anonymous):

so which one would it be?

OpenStudy (aaronq):

so (1)sin(28)=(1.55)(sinx)

OpenStudy (anonymous):

it said error?

OpenStudy (anonymous):

is it 26 degrees?

OpenStudy (aaronq):

i keep getting 17.63

OpenStudy (anonymous):

are you sure thats the right formula?

OpenStudy (aaronq):

yeah I'm pretty sure

OpenStudy (anonymous):

i dont understand

OpenStudy (aaronq):

sin28 = 0.469 0.469/1.55 = sin x sinx= 0.3028 asin(0.3028)=x = 17.63

OpenStudy (anonymous):

but thats none of the anwers

OpenStudy (aaronq):

idk thats what the math is saying :S

OpenStudy (aaronq):

omg I'm sorry i read it backwards

OpenStudy (aaronq):

it works out now

OpenStudy (anonymous):

so what would it be?

OpenStudy (aaronq):

plug it into a calculator, i can't so bluntly give you the answer

OpenStudy (aaronq):

(1.55)(sin28)=(1)(sinx)

OpenStudy (anonymous):

my calculator doesnt work!:/

OpenStudy (aaronq):

http://web2.0calc.com/ :)

OpenStudy (anonymous):

ok thnks

OpenStudy (anonymous):

so what was it?

OpenStudy (shannonmcc10):

Answer is 47 degrees